OpenStudy (anonymous):

please help find domain/range of this function f(x)=-sqrt(x+3)

Vocaloid (vocaloid):

well, let's start off w/ the domain everything under the square root (radical) must be greater than or equal to 0, giving us: \[x+3 \ge0\]

OpenStudy (anonymous):

i understand that... but the negative before the sqrt is what makes me unsureee

Vocaloid (vocaloid):

the negative sign doesn't make a difference for domain. it will change the range though

OpenStudy (anonymous):

then i think its X must be greater than or equal to -3 for domain

Vocaloid (vocaloid):

yup! do you think you can find the range now?

OpenStudy (anonymous):

umm.. no haha.. plug -3 in for x..? :s

Vocaloid (vocaloid):

not quite, haha. well, first, let's consider y = sqrt(x+3), without the negative sign for now you would agree that y must be greater than or equal to 0, correct? since square root of anything must be greater than or equal to 0

OpenStudy (anonymous):

yes id agree

Vocaloid (vocaloid):

all the negative sign does is "flips" the graph across the y-axis, making our domain y \[y \le 0\] since all our y values are now negative or 0

Vocaloid (vocaloid):

*x-axis

Vocaloid (vocaloid):

our graph ends up looking like this: 

Vocaloid (vocaloid):

*range

OpenStudy (anonymous):

so the range would be (-infinity,0)

Vocaloid (vocaloid):

almost, (-infinity,0] since 0 is included

Vocaloid (vocaloid):

(we use square brackets when we want to include the other value)

OpenStudy (anonymous):

ohhhh yaaaa, sorry brain fart lol thank you so much

OpenStudy (anonymous):

@Vocaloid i have another question... what would be the best way to find f(x)=3sinx

Vocaloid (vocaloid):

ah, ok, let's start off by looking at f(x) = sin(x) this function has a domain of (-infinity,infinity) and a range of [-1,1] and looks something like this: